Three individuals form a partnership and agree to divide the profits equally. X invests $9,000, Y invests $7,000, Z invests $4,000. If the profits are $4,800, how much less does x receive compared to having the profits divided in proportion to the amounts invested by X, Y, and Z?
Answer by Theo(13342)

x invests 9000
y invests 7000
z invests 4000
total investment equals $20,000
-----
profits are 4800 / 3 = 1600 apiece
-----
if they each received profits in the same ratio as their investments,
x would have received 9/20 * 4800 = 2160
y would have received 7/20 * 4800 = 1680
z would have received 4/20 * 4800 = 960
-----
total profits to x, y, and z would be 2160 + 1680 + 960 = 4800
-----
Answer to your question is:
x receives 1600
x would have received 2160
x receives 2160 - 1600 = 560 less than x would have received had the profits been divided based on the ratio of investments.
-----